OpenStudy (anonymous):

2/3x+1/9y=2 whats the y-intercept for this linear equation.

OpenStudy (anonymous):

1/9y=-2/3x+2 y=-6x+18 y int =18

OpenStudy (anonymous):

slope intercept form is y=mx+b m=slope bis y intercept